OpenStudy (anonymous):

can anyone please explain to me how to find a polynomial of degree three with the given numbers as zeros? -6,-√ 3,√ 3

OpenStudy (anonymous):

(you do not have to multiply out)

OpenStudy (anonymous):

'the given numbers as zeros? -6,-√ 3,√ 3 ' what do u mean by this?

OpenStudy (anonymous):

are these the roots of the polynomial? -6,-√ 3,√ 3

OpenStudy (anonymous):

yes I believe that's what it means, that's just how it was worded

OpenStudy (anonymous):

(x+6)(x+√ 3)(x-√ 3)

OpenStudy (anonymous):

Just put them in like that. You can multiply them out if you so wish.

OpenStudy (anonymous):

Well, that would be the polynomial if you multiplied it out.

OpenStudy (anonymous):

(x+√ 3)(x-√ 3) x^2-x√ 3+x√ 3-3 (X+6)(X^2-3) X^3-3x+6x^2-18 \[x^3+6x^2-3x-18\]

OpenStudy (anonymous):

@Jacquelinenorth did that kid 'explain' it so you understand...?

OpenStudy (anonymous):

I think so...I'm trying the steps right now by myself

OpenStudy (anonymous):

k

OpenStudy (anonymous):

okay I understand this one but my next one has i and -i in it

OpenStudy (anonymous):

Basically, if a zeros are -5 and 2, it would be (x+5)(x-2) Therefore, it is (x+√ 3)(x-√ 3)(x+6) The first is for the -√3 zero, the second is for the √3 zero, and the third is for the -6 zero. If you multiply these three out, they give you the full equation of the polynomial.

OpenStudy (anonymous):

Oh. Well. Grr. Those would theoretically be graphed in complex coordinate planes, because they involve non-real numbers.

OpenStudy (anonymous):

oh....yeah....all that stuff you said

OpenStudy (anonymous):

Well, anyways, hit me with your best shot. I'll try.

OpenStudy (anonymous):

okay it says: find a polynomial of degree four with the given numbers as zeros: i,-i,-2+√ 5,-2-√ 5

OpenStudy (anonymous):

so I would have (x+i) (x-i) (x-2+√ 5)(x+2-√ 5)?

OpenStudy (anonymous):

I'm not entirely certain that it would be x. Because once we switch over to the other coordinate plane it is now in the (a,b) format, not (x,y) which complicates things.

OpenStudy (anonymous):

oh geez louise

OpenStudy (anonymous):

Actually.

OpenStudy (anonymous):

If \[i=\sqrt-1\] then \[i^2=\sqrt-1 \times \sqrt -1\] \[=-1\]

OpenStudy (anonymous):

Hold on, let me try.

OpenStudy (anonymous):

I think I murdered you with math on accident

OpenStudy (anonymous):

(x+i) (x-i) x^2-xi+xi-sqrt(-1^2) \[x^2-x*i+x*i-\sqrt(-1)^2\] So I was correct. X^2--1 x^2+1=(x+i)(x-i) Now, all we have to do is multiply out the other two factors, which are (x-2+√ 5)(x+2-√ 5) X^2+2x-x√ 5-2x-4+2√ 5+x√ 5+2√ 5-5 \[x^2+2x-2x-x√ 5+x√ 5+2√ 5+2√ 5-5-4\] Now this is really nasty looking. \[x^2-x√ 5+x√ 5+4√ 5-5-4\] Sweet Jesus Combine them nasty like terms now. \[x^2+4√ 5-9\] Oh my effing god I'm not even done yet, where is my gun because I'm dying over here.

OpenStudy (anonymous):

Yes you did I'm dying over here thank you ;)

OpenStudy (anonymous):

I'm sorry!!!! do not die without finishing this equation haha

OpenStudy (anonymous):

\[(x^2+1)(x^2+4√ 5-9)\] x^4+4√ 5x^2-9x^2+x^2+4√ 5-9 \[x^4+(4√ 5)x^2 -8x^2+4√ 5-9\] I don't really even know how to get rid of that √ 5 without having the redo every single bit of this, and quite frankly I don't think I need to do that all over again. That was seriously one of the most mind-intensive things I've ever done. Even though it isn't really anything that new, it took all over my attention and all of my effort just to do that. O.o and it's probably wrong too but at this point it is what it is, I'm dyin' over here.

OpenStudy (anonymous):

@halorazer haha well thanks for trying, I'll double check it with my instructor tomorrow. thanks for being so patient, I won't force you to do any more math. and sorry for the math murder

OpenStudy (anonymous):

I don't have a problem with the math as long as it isn't that. ;o

OpenStudy (anonymous):

deal!